A cosmic ray proton moving toward the Earth at 6.9 × 107 m/s experiences a magnetic force of 1.3 × 10−16 N. What is the strength of the magnetic field (in uT) if there is a 42° angle between it and the proton’s velocity?


answer to 2 decimal places


1
Expert's answer
2021-05-26T12:54:26-0400

Gives

Velocity (v) "=6.9\times10^7m/sec"

Force"(F)=1.3\times10^{-16}N"

"\theta=42°"

We know that

Magnetic force

"F=q(v\times B)"

"F=qvBsin\theta"

"B=\frac{F}{qvsin42°}"

"B=\frac{1.3\times10^{-16}}{1.6\times10^{-19}\times6.9\times10^{7}\times sin42°}"

B="1.76\times10^{-5}T"

B"=17.6\times10^{-6} T"

B="17.6\mu T"